# New gravitational wave polarization modes in the torsionless spacetime

### v2

@@ I Introduction
We adopt the convention $c=G=1$ and use the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ). The indices $(\mu,\nu,\lambda,\rho)$ correspond to four-dimensional spacetime indices, ranging from 00 to $3$, while the indices $(i,j,k,l)$ represent three-dimensional spatial indices, ranging from $1$ to $3$, corresponding to the $(+x,+y,+z)$ directions, respectively. $\Gamma^{\lambda}_{~{}\mu\nu}$ denotes a general connection, while $\widehat{\Gamma}^{\lambda}_{~{}\mu\nu}$ refers to the Levi-Civita connection. Similarly, for any quantity $A$ that depends on the connection, $\widehat{A}=A\arrowvert_{\Gamma^{\lambda}_{~{}\mu\nu}=\widehat{\Gamma}^{\lambda}_{~{}\mu\nu}}$.
@@ II Gauge invariants in the linearized Palatini theory
Then, we can take the background perturbations as
@@ III.2 The equation of motion in a torsionless spacetime
^$\widehat{\nabla}_{\nu}\mathcal{H}_{\lambda}^{~{}\mu\nu}\!$ $=$ $\!2\left(\Sigma^{\mu}_{~{}\lambda}-T^{\mu}_{~{}\lambda}\right)-N^{\mu}_{~{}\alpha\beta}\mathcal{H}_{\lambda}^{~{}\alpha\beta}+N^{\alpha}_{~{}\lambda\beta}\mathcal{H}_{\alpha}^{~{}\mu\beta},$ (25)
^$\widehat{\nabla}_{\mu}\Sigma^{\mu}_{~{}\alpha}\!$ $=$ $\!\frac{1}{4}\widehat{R}_{\lambda\mu\nu\alpha}\left(\mathcal{H}^{\lambda\mu\nu}-\mathcal{H}^{\mu\lambda\nu}\right)+\frac{1}{2}\widehat{\nabla}_{\lambda}\left(\mathcal{H}^{\mu\nu\lambda}N_{\mu\nu\alpha}\right)-\frac{1}{2}\mathcal{H}_{\lambda}^{~{}\mu\nu}\widehat{\nabla}_{\alpha}N^{\lambda}_{~{}\mu\nu}.$ (26)
@@
In fact, the above assumption can be relaxed without changing anything. A proof of this is provided in Appendix B.
@@
By multiplying both sides of the equation by $u_{\mu}$ and using $u_{\mu}u^{\mu}=-1$ and $u_{\mu}\widehat{D}u^{\mu}/d\tau=0$, it can be concluded that
@@
From Eq. (36), it can be seen that the motion of a free particle is only related to its initial position and velocity, and is independent of the specific values of its charges $m$, $H_{\mu}$, and $P_{\mu}$. This is an interesting property because it suggests that we can still interpret the free motion of particles as geometric effects in this case. For a general metric-affine spacetime, the motion of particles is often related to the specific values of their charges Damianos Iosifidis .
In the torsionless spacetime we are considering, $N^{\lambda}_{~{}\mu\nu}$ depends solely on the non-metricity $Q_{\lambda\mu\nu}\coloneqq\nabla_{\lambda}g_{\mu\nu}$, specifically, $N^{\lambda}_{~{}\mu\nu}=\frac{1}{2}Q^{\lambda}_{~{}\mu\nu}-Q_{(\mu~{}\nu)}^{~{}~{}\lambda}$ Lavinia Heisenberg2 . Therefore, whether a particle possesses a hypermomentum charge $H_{\mu}$ determines whether it can experience the effects of non-metricity. If a particle does not carry a hypermomentum charge $H_{\mu}$ or if spacetime is free from non-metricity ($N^{\lambda}_{~{}\mu\nu}=0$), the equation of motion of particles will reduce to Eq. (18), which is identical to the equation in general relativity. It is important to note that the motion of particles exhibits a discontinuous variation in this context. In spacetime with non-zero non-metricity, a free particle carrying hypermomentum will be affected by non-metricity, regardless of how small its parameter $H_{\mu}$ is, as long as $H_{\mu}\neq 0$. In this case, the particle’s motion behaves identically to that of a particle with large hypermomentum charge, given the same initial position and velocity. This is in contrast to the motion of a particle that does not carry hypermomentum, which follows a different trajectory.
@@ IV Shear mode: a new class of gravitational wave polarization modes
$ds^{2}\approx\!$displaystyle-- $\!dt^{2}\left[1+2a_{i}\delta x^{i}+\left(a_{i}\delta x^{i}\right)^{2}+\widehat{R}_{i0j0}\delta x^{i}\delta x^{j}\right]$ (39)
@@
[…]
@@
Let us first review the case where the equation of motion satisfies (18), which indicates that the test particles do not carry a hypermomentum charge. At this stage, the equation of relative motion for particles is exactly what we know as landau ; MTW ; Michele Maggiore
@@
We consider plane gravitational waves propagating in the $+z$ direction, without loss of generality. This means that
@@
Figure 1: Six polarization modes of gravitational waves Eardley . The gravitational waves propagate in the $+z$ direction. The solid line represents the shape of the test particle array—initially spherical—depicting the motion of particles relative to the central particle of the sphere when the wave phase is $\frac{\pi}{2}$. The dotted line illustrates the shape of the array at a phase of $\frac{3}{2}\pi$. No relative motion occurs between test particles along the third axis, which is not shown in the figure.
@@
Since $P_{4}$ and $P_{5}$ are induced by spatial tensors, we classify them as tensor modes. Similarly, $P_{2}$ and $P_{2}$ correspond to vector modes, while $P_{1}$ and $P_{6}$ are scalar modes.
@@
it becomes apparent that the additional term $\partial_{j}N_{i00}$ breaks the symmetry of $\left(i,j\right)$. Therefore, the number of independent components in $A_{ij}$ exceeds six, implying the existence of more than six independent polarization modes of gravitational waves. Given that the gravitational waves we are considering propagate in the $+z$ direction, only $\partial_{3}N_{i00}$ in $\partial_{j}N_{i00}$ can be non-zero. So, we can define the independent polarization modes of gravitational waves as
@@
We use $P_{7}$ as an example to illustrate how the test particles move in this new mode. Note that by rotating the $x-y$ plane $90$ degrees, $P_{7}$ is transformed into $P_{8}$, so the motion modes of $P_{8}$ and $P_{7}$ are identical. Given that only $P_{7}\neq 0$ among all polarization modes, substituting Eq. (60) into Eq. (47) yields
@@
It is evident that a $P_{7}$-mode gravitational wave propagating along the $+z$ direction does not induce relative changes in the $y$ and $z$ directions of the test particles. However, it will cause periodic relative motion in the $x$ direction for test particles with an initial relative separation in the $z$ direction. This motion mode is tangential, so we refer to $P_{7}$ and $P_{8}$ as shear modes, specifically the shear-$x$ mode and shear-$y$ mode, respectively. In Fig. 2 and Fig. 3, we show the motion of the test particles in these two new modes.
Figure 2: The two new shear modes. The gravitational waves in the figure propagate in the $+z$ direction. In the shear-$x$ and shear-$y$ modes, the periodic tangential motion of a test particle array occurs along the $x$ and $y$ directions, respectively. In addition, the amplitude of this motion increases with the relative distance in the $z$-direction between the test particles. The solid and dashed lines represent the cases where the phases are labeled as $\frac{\pi}{2}$ and $\frac{3}{2}\pi$, respectively.
@@
The analysis of gravitational wave polarization modes in this theory requires solving the linearized field equations. As shown in Appendix D and Refs. Y.Dong5 ; Y.Dong4 ; Y.Dong3 , these equations can be simplified by decoupling them using gauge invariants and transforming them into algebraic equations in momentum space, thereby rendering the system more tractable. By applying these methods, we find that for tensor modes, the $+$ and $\times$ polarizations always propagate at the speed of light, regardless of whether the particle carries hypermomentum.
@@
For this solution, the corresponding gauge invariant vector perturbations satisfy
@@
Since $\Xi_{i}=0$, no vector polarization mode can be detected for test particles without hypermomentum charge, as indicated by Eq. (58). However, for test particles with hypermomentum charges, Eq. (IV) reveals the existence of shear-$x$ and shear-$y$ modes with mass $m$ that satisfy Eq. (65) at this time.
@@
From Eq. (58), it follows that, due to $\Theta=\phi=0$, test particles without hypermomentum charge cannot detect any scalar modes. In contrast, according to Eq. (IV), test particles with a hypermomentum charge will detect a longitudinal mode with mass $m$ satisfying Eq. (67).
In our simple example, test particles without hypermomentum charge can only detect tensor modes propagating at the speed of light, similar to general relativity. However, for particles with a hypermomentum charge, additional shear and longitudinal modes with mass $m^{2}=\alpha/\beta$ may also be detected.
@@ VI Conclusion
We first examined the action of the matter field $S_{m}$ and derived the equation of motion for free particles in a torsionless spacetime using generalized covariance. In a torsionless spacetime, free particles without a hypermomentum charge follow the same equation of motion (18) as in general relativity. However, for particles carrying hypermomentum charges, the equation of motion must be modified (see Eq. (36)). In this modified equation, the particle’s charges $m$, $P^{\mu}$, and $H^{\mu}$ do not appear. Therefore, the trajectory of the particles depends only on their initial position and velocity, and is independent of the specific values of the charges they carry. From this perspective, the effect of a hypermomentum charge is solely to introduce non-metricity interactions in spacetime, while the free motion of particles remains a manifestation of geometric effects. This is particularly interesting because, in a general metric-affine spacetime, the trajectory of particles would depend on the specific values of the charges they carry Damianos Iosifidis . It is also important to note that the equation of motion for a free particle does not correspond to the geodesic equation in a metric-affine geometry.
@@
Having identified, from the perspective of the equations of motion, that non-metricity induces two novel gravitational wave polarization modes, we proceeded to examine the field equations of the theory. Through a concrete example, we demonstrated that these new polarization modes can indeed exist in certain theories. In the example theory we considered, test particles without a hypermomentum charge can only detect the $+$ and $\times$ modes, which propagate at the speed of light, as in general relativity. However, for particles with a hypermomentum charge, shear and longitudinal modes with mass $m^{2}=\alpha/\beta$ can also be detected. This example highlights two important aspects of gravitational wave polarization in a torsionless spacetime. First, the polarization effect differs between particles that carry hypermomentum and those that do not. Second, it suggests that new shear modes may indeed exist.

### v3

@@ I Introduction
[…]
@@ II Gauge invariants in the linearized Palatini theory
At this background value, the theory maintains Lorentz and CPT symmetries. Then, we can take the background perturbations as
@@ III.2 The equation of motion in a torsionless spacetime
In the current theoretical framework of matter fields (i.e., the Standard Model), there is no direct coupling between matter fields and the connection. Nevertheless, one can hypothetically introduce such a coupling without contradicting existing solar-system experiments. As an illustrative example, consider a scalar matter field $\phi$ governed by the action
$S_{m}=\int d^{4}{x}\sqrt{-g}\left[-\frac{1}{2}\partial_{\mu}\phi\!~\!\partial^{\mu}\phi-\frac{1}{2}m^{2}\phi^{2}-\xi R\phi^{2}\right].$ (20)
Here, $R\coloneqq g^{\nu\sigma}R^{\mu}_{~\nu\mu\sigma}$, and $R^{\mu}_{~\nu\lambda\sigma}$ is the curvature tensor. In the solar system, where $R$ is extremely small, the parameter $\xi$ can be tuned so that the dynamics of $\phi$ are indistinguishable from those of a Klein-Gordon field with mass $m$. In contrast, in strong gravitational fields, $\phi$ exhibits behavior that deviates from the standard Klein-Gordon field. Moreover, as noted in Ref. Damianos Iosifidis6 , introducing a coupling between matter fields and the connection provides new avenues for exploring questions related to cosmic evolution.
@@
$\widehat{\nabla}_{\nu}\mathcal{H}_{\lambda}^{~\mu\nu}\!$ $=$ $\!2\left(\Sigma^{\mu}_{~\lambda}-T^{\mu}_{~\lambda}\right)-N^{\mu}_{~\alpha\beta}\mathcal{H}_{\lambda}^{~\alpha\beta}+N^{\alpha}_{~\lambda\beta}\mathcal{H}_{\alpha}^{~\mu\beta},$ (26)
$\widehat{\nabla}_{\mu}\Sigma^{\mu}_{~\alpha}\!$ $=$ $\!\frac{1}{4}\widehat{R}_{\lambda\mu\nu\alpha}\left(\mathcal{H}^{\lambda\mu\nu}-\mathcal{H}^{\mu\lambda\nu}\right)+\frac{1}{2}\widehat{\nabla}_{\lambda}\left(\mathcal{H}^{\mu\nu\lambda}N_{\mu\nu\alpha}\right)-\frac{1}{2}\mathcal{H}_{\lambda}^{~\mu\nu}\widehat{\nabla}_{\alpha}N^{\lambda}_{~\mu\nu}.$ (27)
@@
In fact, the above assumption can be relaxed without changing anything. A proof of this is provided in Appendix B. It is worth noting that, in Appendix B, we demonstrate that assumptions (29)–(31) are equivalent to assumptions (82)–(84), which correspond to considering only the monopole moment of the body. Therefore, our assumptions actually correspond to the monopole approximation in the multipole expansion.
@@
This corresponds to the torsionless spacetime case of Eq. (99) in Ref. Dirk Puetzfeld1 . By multiplying both sides of the equation by $u_{\mu}$ and using $u_{\mu}u^{\mu}=-1$ and $u_{\mu}\widehat{D}u^{\mu}/d\tau=0$, it can be concluded that
@@
From Eq. (37), it can be seen that the motion of a free particle is only related to its initial position and velocity, and is independent of the specific values of its charges $m$, $H_{\mu}$, and $P_{\mu}$. This is an interesting property because it suggests that we can still interpret the free motion of particles as geometric effects in this case. One might argue that the independence of motion from the magnitude of the hypermomentum implies that hypermomentum cannot be measured, which could appear physically counterintuitive. However, in our analysis, we consider only the gravitational fields (the metric and the connection) and neglect any external fields. In addition, we work within the monopole approximation. It is therefore plausible that, when external fields are included or higher-order multipole moments are considered, the magnitude of the hypermomentum would influence the motion of the body, rendering the hypermomentum, in principle, measurable. For a general metric-affine spacetime, the motion of particles is often related to the specific values of their charges Damianos Iosifidis . Finally, we note that the validity of the monopole approximation requires that, in the equations under consideration, the contributions from hypermomentum and the energy-momentum tensor are of the same order of magnitude. In situations where the hypermomentum contribution is small compared to that of the energy-momentum tensor, it can be neglected at leading order, reducing the equations of motion to Eq. (18).
In the torsionless spacetime we are considering, $N^{\lambda}_{~\mu\nu}$ depends solely on the non-metricity $Q_{\lambda\mu\nu}\coloneqq\nabla_{\lambda}g_{\mu\nu}$, specifically, $N^{\lambda}_{~\mu\nu}=\frac{1}{2}Q^{\lambda}_{~\mu\nu}-Q_{(\mu~\nu)}^{~~\lambda}$ Lavinia Heisenberg2 . Therefore, whether a particle possesses a non-negligible hypermomentum charge $H_{\mu}$ determines whether it can experience the effects of non-metricity. If a particle does not carry a hypermomentum charge $H_{\mu}$ (or if the hypermomentum contribution is negligible) or if spacetime is free from non-metricity ($N^{\lambda}_{~\mu\nu}=0$), the equation of motion of particles will reduce to Eq. (18), which is identical to the equation in general relativity.
@@ IV Shear mode: a new class of gravitational wave polarization modes
$ds^{2}\approx\!$ - $\!dt^{2}\left[1+2a_{i}\delta x^{i}+\left(a_{i}\delta x^{i}\right)^{2}+\widehat{R}_{i0j0}\delta x^{i}\delta x^{j}\right]$ (40)
@@
$\eta^{i}\left(t\right)=\bar{\eta}^{i}-\bar{\eta}^{j}\int_{0}^{t}dt^{\prime}\int_{0}^{t^{\prime}}dt^{\prime\prime}A^{i}_{~j}\left(t^{\prime\prime}\right).$ (50)
@@
Let us first review the case where the equation of motion satisfies (18), which indicates that the test particles do not carry a hypermomentum charge (or the effect of hypermomentum is negligible). At this stage, the equation of relative motion for particles is exactly what we know as landau ; MTW ; Michele Maggiore
@@
We consider plane gravitational waves propagating in the +z+z direction, without loss of generality. This means that
@@
Figure 1: Six polarization modes of gravitational waves Eardley . The gravitational waves propagate in the +z+z direction. The solid line represents the shape of the test particle array—initially spherical—depicting the motion of particles relative to the central particle of the sphere when the wave phase is $\frac{\pi}{2}$. The dotted line illustrates the shape of the array at a phase of $\frac{3}{2}\pi$. No relative motion occurs between test particles along the third axis, which is not shown in the figure.
@@
Since $P_{4}$ and $P_{5}$ are induced by spatial tensors, we classify them as tensor modes. Similarly, $P_{2}$ and $P_{3}$ correspond to vector modes, while $P_{1}$ and $P_{6}$ are scalar modes.
@@
it becomes apparent that the additional term $\partial_{j}N_{i00}$ breaks the symmetry of $\left(i,j\right)$. Therefore, the number of independent components in $A_{ij}$ exceeds six, implying the existence of more than six independent polarization modes of gravitational waves. Given that the gravitational waves we are considering propagate in the +z+z direction, only $\partial_{3}N_{i00}$ in $\partial_{j}N_{i00}$ can be non-zero. So, we can define the independent polarization modes of gravitational waves as
@@
We use $P_{7}$ as an example to illustrate how the test particles move in this new mode. Note that by rotating the xyx-y plane $90$ degrees, $P_{7}$ is transformed into $P_{8}$, so the motion modes of $P_{8}$ and $P_{7}$ are identical. Given that only $P_{7}\neq 0$ among all polarization modes, substituting Eq. (61) into Eq. (48) yields
@@
It is evident that a $P_{7}$-mode gravitational wave propagating along the +z+z direction does not induce relative changes in the $y$ and $z$ directions of the test particles. However, it will cause periodic relative motion in the $x$ direction for test particles with an initial relative separation in the $z$ direction. This motion mode is tangential, so we refer to $P_{7}$ and $P_{8}$ as shear modes, specifically the shear-$x$ mode and shear-$y$ mode, respectively. In Fig. 2 and Fig. 3, we show the motion of the test particles in these two new modes.
Figure 2: The two new shear modes. The gravitational waves in the figure propagate in the +z+z direction. In the shear-$x$ and shear-$y$ modes, the periodic tangential motion of a test particle array occurs along the $x$ and $y$ directions, respectively. In addition, the amplitude of this motion increases with the relative distance in the $z$-direction between the test particles. The solid and dashed lines represent the cases where the phases are labeled as $\frac{\pi}{2}$ and $\frac{3}{2}\pi$, respectively.
@@
It should be noted that polarization modes of gravitational waves and the propagating degrees of freedom of perturbations are distinct concepts. Polarization modes are defined via the relative motions of test particles and therefore represent modes of motion. From this perspective, the number of independent polarization modes in a theory does not directly determine the number of propagating degrees of freedom of the perturbations. The presence of a given polarization mode only indicates that the corresponding perturbation component can propagate; to determine the actual number of propagating degrees of freedom in a theory, one must solve the field equations.
@@
The analysis of gravitational wave polarization modes in this theory requires solving the linearized field equations. As shown in Appendix D and Refs. Y.Dong5 ; Y.Dong4 ; Y.Dong3 , these equations can be simplified by decoupling them using gauge invariants and transforming them into algebraic equations in momentum space, thereby rendering the system more tractable. By applying these methods, we find that for tensor modes, the ++ and $\times$ polarizations always propagate at the speed of light, regardless of whether the particle carries hypermomentum.
@@
To avoid the emergence of an imaginary mass, we require that $\alpha$ and $\beta$ share the same sign. For this solution, the corresponding gauge invariant vector perturbations satisfy
@@
Since $\Xi_{i}=0$, no vector polarization mode can be detected for test particles without hypermomentum charge (or the effect of hypermomentum is negligible), as indicated by Eq. (59). However, for test particles with hypermomentum charges, Eq. (IV) reveals the existence of shear-$x$ and shear-$y$ modes with mass $m$ that satisfy Eq. (66) at this time.
@@
From Eq. (59), it follows that, due to $\Theta=\phi=0$, test particles without hypermomentum charge (or the effect of hypermomentum is negligible) cannot detect any scalar modes. In contrast, according to Eq. (IV), test particles with a hypermomentum charge will detect a longitudinal mode with mass $m$ satisfying Eq. (68).
In our simple example, test particles without hypermomentum charge (or the effect of hypermomentum is negligible) can only detect tensor modes propagating at the speed of light, similar to general relativity. However, for particles with a hypermomentum charge, additional shear and longitudinal modes with mass $m^{2}=\alpha/\beta$ may also be detected.
@@ VI Conclusion
We first examined the action of the matter field $S_{m}$ and derived the equation of motion for free particles in a torsionless spacetime using generalized covariance. In a torsionless spacetime, free particles without a hypermomentum charge (or the effect of hypermomentum is negligible) follow the same equation of motion (18) as in general relativity. However, for particles carrying hypermomentum charges, the equation of motion must be modified (see Eq. (37)). In this modified equation, the particle’s charges $m$, $P^{\mu}$, and $H^{\mu}$ do not appear. Therefore, the trajectory of the particles depends only on their initial position and velocity, and is independent of the specific values of the charges they carry. From this perspective, the effect of a hypermomentum charge is solely to introduce non-metricity interactions in spacetime, while the free motion of particles remains a manifestation of geometric effects. This is particularly interesting because, in a general metric-affine spacetime, the trajectory of particles would depend on the specific values of the charges they carry Damianos Iosifidis . It is also important to note that the equation of motion for a free particle does not correspond to the geodesic equation in a metric-affine geometry.
@@
Having identified, from the perspective of the equations of motion, that non-metricity induces two novel gravitational wave polarization modes, we proceeded to examine the field equations of the theory. Through a concrete example, we demonstrated that these new polarization modes can indeed exist in certain theories. In the example theory we considered, test particles without a hypermomentum charge can only detect the ++ and $\times$ modes, which propagate at the speed of light, as in general relativity. However, for particles with a hypermomentum charge, shear and longitudinal modes with mass $m^{2}=\alpha/\beta$ can also be detected. This example highlights two important aspects of gravitational wave polarization in a torsionless spacetime. First, the polarization effect differs between particles that carry hypermomentum and those that do not. Second, it suggests that new shear modes may indeed exist.
@@
Finally, it is necessary to discuss the relationship between nonmetricity and Lorentz invariance. Although nonmetricity is often presumed to be subject to strict constraints from Lorentz symmetry, the two concepts are not intrinsically or logically connected. This belief may stem from two main viewpoints. One possible viewpoint is that, similar to vector-tensor theories, assigning a nonzero background value to nonmetricity can introduce a preferred direction in spacetime. As a result, fields propagating on such a background may exhibit Lorentz symmetry breaking, which could further lead to CPT violation. However, in our work, the background field is defined as in Eq. (3); that is, the background value of nonmetricity is taken to be zero. Therefore, Lorentz invariance is preserved, and CPT symmetry is consequently maintained. This is also evident from the most general second-order perturbation action given in Eq. (98). Another viewpoint holds that nonmetricity modifies the inner product of vectors under parallel transport, allowing a null vector to become time-like and thereby distorting the light cone structure. This argument, however, rests on the assumption that light propagation is governed by the parallel transport of its wave vector—an assumption valid in general relativity but not necessarily applicable to modified gravity theories. In fact, parallel transport is a mathematical notion, and the physical trajectory of light must be derived from the electromagnetic action in a given background through the field equations and geometric optics approximation. Therefore, the existence of nonmetricity does not inherently imply a breakdown of the light cone structure, and consequently, does not necessarily lead to Lorentz symmetry violation. Notably, as demonstrated in Ref. Lavinia Heisenberg1 , it is possible to construct a theory involving nonmetricity that remains physically equivalent to general relativity. Based on the above discussion, the result of this paper are consistent with the constraints imposed by current experiments S. Herrmann ; S. Navas .